OpenStudy (anonymous):

5x + 3 = 2x + 15 x =

OpenStudy (imtant):

Same goes to the previous one, 5x-2x=15-3 3x=12 so x=?

OpenStudy (anonymous):

4!!!!!

OpenStudy (imtant):

Yes..